# The externally definable Ramsey property and fixed points on type spaces

### v3

@@ 7.1. The infinite lexicographic product
The following definition is a special case of [22, Example 4.6].
@@
Let $\left(M_{i}\right)_{i\in\omega}$ be a family of countably infinite $\mathcal{L}$-structures, where for all $i\in\omega$ the universe of $M_{i}$ is $\omega$. The infinite lexicographic product $\prod_{\omega^{<\omega}}M_{i}$ is the $\mathcal{L}$-structure with universe $\omega^{\omega}$ where, for every $a_{1},\dots,a_{k}\in\omega^{\omega}$ and every $k$-ary relation $R\in\mathcal{L}$, letting $n\in\omega$ be maximal such that $a_{1}\upharpoonright n=\dots=a_{k}\upharpoonright n$, we have that
$\prod_{\omega^{<\omega}}M_{i}\models R(a_{1},\dots,a_{k})\iff M_{n}\models R(a_{1}(n+1),\dots,R(a_{k}(n+1)).$
While most of the results in the following can be stated more generally (see [22, Definition 4.5]), for our purposes, we restrict our attention to products as in Definition 7.9.
Furthermore, throughout this section, we fix a family $\left(M_{i}\right)_{i\in\omega}$ of countable structures, where each $M_{i}$ is transitive, ultrahomogeneous and has universe $\omega$. (This is called $\omega$-pure in [22, Definition 5.2].)
We fix a family $\Set{s_{i}}{i<\omega}$ of binary relation symbols, and let $M_{i}^{s}$ be the expansion of $M_{i}$ by $\Set{s_{n}}{n<\omega}$ such that $\left(s_{n}\right)^{M_{i}^{s}}=\left\{\begin{array}[]{ll}\left(M_{i}\right)^{2}&\mbox{if }i<n\\ \varnothing&\mbox{if }i\geq n.\end{array}\right.$.
Remark 7.10 ([22, Remark 5.4]).
With the expansion above, $\prod_{\omega^{<\omega}}M_{i}^{s}$ has refining equivalence relations $s_{0}\supsetneq s_{1}\supsetneq s_{2}\supsetneq\cdots$. We also have that $\prod_{\omega^{<\omega}}M_{i}^{s}$ is the expansion of $\prod_{\omega^{<\omega}}M_{i}$ by $\Set{s_{i}}{i<\omega}$ such that $s_{i}(x,y)\Leftrightarrow\operatorname{height}(x\wedge y)>i$.
The following lemma is a special case of [22, Lemma 4.15]. For completeness of presentation, we present an elementary proof here.
@@
Let $N_{i}:=M_{i+1}$ and let $N^{s}:=\prod_{\omega^{<\omega}}N_{i}^{s}$. Then we have $\prod_{\omega^{<\omega}}M_{i}^{s}\cong M_{0}[N]^{s_{0}}$.
Proof.
Let $g:M_{0}[N^{s}]^{s_{0}}\to\prod_{\omega^{<\omega}}M_{i}^{s}$ be defined as $g((a,b)):=a\smallfrown~{}b$ (the concatenation of $a$ and $b$). Then $g$ is clearly surjective. To show $g$ is an isomorphism:
  1. (1)
    For $(a,b),(c,d)\in M_{0}[N^{s}]^{s_{0}}$:
    $M_{0}[N^{s}]^{s_{0}}\models s_{0}((a,b),(c,d))$ $\iff a=c$
    $\iff g((a,b))(0)=g((c,d))(0)$
    $\iff M_{i}^{s}\models s_{0}(g((a,b)),g((c,d))).$
  2. (2)
    For every $n\in\omega\setminus\Set{0}$ and every $(a,b),(c,d)\in M_{0}[N^{s}]^{s_{0}}$:
    $M_{0}[N^{s}]^{s_{0}}\models s_{i}((a,b),(c,d))$
    $\iff a=c\land N^{s}\models s_{n}(b,d)$
    $\iff a=c\land\operatorname{height}((0)\smallfrown b\wedge(0)\smallfrown d)\geq n$
    $\iff\operatorname{height}((a\smallfrown b)\wedge(b\smallfrown d))\geq n$
    $\iff\prod_{\omega^{<\omega}}M_{i}^{s}\models s_{n}(g((a,b)),g((c,d))).$
  3. (3)
    For every $k$-ary $R\in L$ and $(a,b_{1}),\dots,(a,b_{k})\in M_{0}[N^{s}]^{s_{0}}$: let $h:=\operatorname{height}(b_{1}\wedge\dots\wedge b_{k})$. Then
    $M_{0}[N^{s}]^{s_{0}}\models R((a,b_{1}),\dots,(a,b_{n}))$
    $\iff N^{s}\models R(b_{1},\dots,b_{k})$
    $\iff N_{h}\models R(b_{1}(h+1),\dots,b_{k}(h+1))$
    $\iff M_{h+1}\models R(b_{1}(h+1),\dots,b_{k}(h+1))$
    $\iff M_{h+1}\models R((a\smallfrown b_{1})(h+1),\dots,(a\smallfrown b_{k})(h+1))$
    $\iff M_{h+1}\models R(g(a,b_{1})(h+1),\dots,g(a,b_{k})(h+1))$
    $\iff\prod_{\omega^{<\omega}}M_{i}\models R(g(a,b_{1}),\dots,g(a,b_{k})).$
  4. (4)
    For every $k$-ary $R\in L$ and $(a_{1},b_{1}),\dots,(a_{k},b_{k})\in M_{0}[N^{s}]^{s_{0}}$, such that $a_{i}\neq a_{j}$ for some $1\leq i,j\leq k$:
    $M_{0}[N^{s}]^{s_{0}}\models R((a_{1},b_{1}),\dots,(a_{k},b_{n}))$ $\Leftrightarrow M_{0}\models R(a_{1},\dots,a_{k})$
    $\Leftrightarrow\prod_{\omega^{<\omega}}M_{i}\models R(g(a_{1},b_{1}),\dots,g(a_{k},b_{k})).$
Notice also that in both structures $x=y\iff\bigwedge_{i<\omega}s_{i}(x,y)$. So we also have injectivity, and $g$ is an isomorphism. ∎
@@
A substructure $D\subseteq\prod_{\omega^{<\omega}}M^{s}_{i}$ is dense if for all $t\in\omega^{<\omega}$, there is some $d\in D$ such that $d\supsetneq t$.
Fact 7.13 ([22, Theorem 5.14]).
@@
    Up to isomorphism, there is a unique countable dense substructure $D\subseteq\prod_{\omega^{<\omega}}M_{i}^{s}$.
@@
    Such a $D$ is transitive and homogeneous.
Fact 7.14 ([22, Theorem 6.1]).
If $D\subseteq\prod_{\omega^{<\omega}}M_{i}^{s}$ is a countable dense substructure, then $D$ is indivisible.
@@
Let $D\subseteq\prod_{\omega^{<\omega}}M_{i}^{s}$ be a countable dense substructure. Then the following are equivalent:
@@
Let $D\subseteq\prod_{\omega^{<\omega}}M_{i}^{s}$ and let $D^{\prime}\subseteq\prod_{\omega^{<\omega}}N_{i}^{s}$ be countable dense substructures. Then $D\cong M_{0}[D^{\prime}]^{s_{0}}$.
@@
Notice that $M_{0}[D^{\prime}]^{s_{0}}$ is countable and dense in $M_{0}[N^{s}]^{s_{0}}$. By Lemmas 7.11 and 7.13, $D\cong M_{0}[D^{\prime}]^{s_{0}}$. ∎
@@
Let $D\subseteq\prod_{\omega^{<\omega}}M_{i}^{s}$ be a countable dense substructure. Then $D$ has FPTn iff $M_{0}$ has FPTn.
@@
Let $D^{\prime}\subseteq\prod_{\omega^{<\omega}}N_{i}^{s}$ be countable and dense. By Lemma 7.16, $D\cong M_{0}[D^{\prime}]^{s_{0}}$. By Proposition 7.8, $M_{0}[D^{\prime}]^{s_{0}}$ has FPTn if and only if $M_{0}$ has FPTn. ∎
@@
By Proposition 7.17, $D$ has FPT if and only if $M_{0}$ has FPT. By Fact 7.13, $D$ is ultrahomogeneous. Recall that $M_{0}$ was also chosen to be ultrahomogeneous and countable. Therefore, by Theorem 5.1, $D$ has the externally definable Ramsey property if and only if $D$ has FPT, and $M_{0}$ has the externally definable Ramsey property if and only if $M_{0}$ has FPT. ∎
@@
If $M_{i}$ is an infinite set in the language of pure equality, for all $i\in\omega$, then a countable dense substructure $D\subseteq\prod_{\omega^{<\omega}}M_{i}^{s}$ is the structure with infinitely many refining equivalence relations (unsaturated version) defined in Section 6.8.
@@
The partial type $\Set{x\neq y}\cup\Set{s_{i}(x,y)}{i\in\omega}$ is finitely satisfiable but not satisfiable in $D$. ∎
@@
Let $\mathcal{L}$ be a countable relational language, and let $T$ be a universal $\mathcal{L}$-theory closed under lexicographic products, i.e., if $M,N\models T$ then $M[N]\models T$. Let $T^{s}$ be the $\mathcal{L}\cup\Set{s_{i}}{i\in\omega}$-theory defined as $T$ together with the $\Set{s_{i}}{i\in\omega}$-theory of infinitely many refining equivalence relations with infinite classes.
@@
Take $A_{0},A_{1}$ to be two distinct countable transitive ultrahomogeneous models of $T$, such that $A_{0}$ has the externally definable Ramsey property. For every $a\in 2^{\omega\setminus\{0\}}$, define $M_{i}(a):=A_{a(i)}$ for $i\in\omega\setminus\{0\}$ and $M_{0}(a):=A$. Let $M(a)$ be a countable dense substructure of $\prod_{\omega^{<\omega}}M_{i}(a)^{s}$. By Fact 7.13, $M(a)$ is ultrahomogeneous. By Theorem 7.15, $M(a)$ has the externally definable Ramsey property. By Proposition 7.19, $M(a)$ is not $\aleph_{0}$-saturated.
It remains to show that $M(a)\not\cong M(b)$ for $a\neq b\in 2^{\omega\setminus\{0\}}$. Assume for a contradiction that $f:M(a)\to M(b)$ is an isomorphism. Let $m\in\omega$ be minimal such that $a(m+1)\neq b(m+1)$. By density of $M(a)$, there are $\Set{\sigma_{i}}_{i\in\omega}\in M(a)$ such that $\sigma_{i}\supsetneq(a\upharpoonright m)\smallfrown(i)$. Notice that $\Set{\sigma_{i}}_{i\in\omega}\in M(a)$ are pairwise $s_{m}$-equivalent, but not $s_{m+1}$-equivalent, and, in addition, they $s_{m+1}$-cover their equivalence class, i.e., for every $\sigma^{\prime}\in M(a)$ such that $M(a)\models s_{m}(\sigma^{\prime},\sigma_{0})$, there is some $i\in\omega$ such that $M(a)\models s_{m+1}(\sigma^{\prime},\sigma_{i})$. Since $g$ is an $\mathcal{L}\cup\Set{s_{i}}{i\in\omega}$-isomorphism, this holds for $\Set{g(\sigma_{i})}_{i\in\omega}$ in $M(b)$ as well. Finally, notice that, by the definition of the infinite lexicographic product, the substructure induced on $\Set{\sigma_{i}}_{i\in\omega}$ in $M(a)$ is isomorphic to $M_{m}(a)$ and the substructure induced on $\Set{g(\sigma_{i})}_{i\in\omega}$ in $M(b)$ is isomorphic to $M_{m}(b)$. But $M_{m}(b)\cong A_{b(m)}\not\cong A_{a(m)}\cong M_{m}(a)$. ∎
@@
  • The class of $n$-hypergraphs, for $n\geq 2$.
  • The class of $K_{m}$-free $n$-hypergraphs for $m\geq n\geq 3$.
  • The class of partially ordered sets.
@@ 8. Further questions
Does there exist an ultrahomogeneous structure $M$ with uncountable age such that $M$ has the externally definable Ramsey property but $\operatorname{Aut}(M)$ does not have FPT?
@@
In this paper, for most examples given of Fraïssé structures $M$ with the externally definable Ramsey property, we show this property by checking $\bar{a}$-Ramsey for all colourings, for $|\bar{a}|<t$ (where $t$ is the maximum arity of $\mathcal{L}$ and we have QE). Can we find interesting examples of Fraïssé structures $M$ where $M$ has the externally definable Ramsey property, but does not have $\bar{a}$-Ramsey for $|\bar{a}|<t$? (See Section 6.2.)

### v4

@@ 7.1. The infinite lexicographic product
The following definition is a special case of [21, Example 4.6].
@@
Let $\left(M_{i}\right)_{i\in\omega}$ be a family of countably infinite $\mathcal{L}$-structures, each with domain $\omega$. The infinite lexicographic product $[M_{i}]_{i<\omega}$ is the $\mathcal{L}$-structure with domain $\omega^{\omega}$ where, for every $a_{1},\dots,a_{k}\in\omega^{\omega}$ and every $k$-ary relation $R\in\mathcal{L}$, letting $n\in\omega$ be maximal such that $a_{1}\upharpoonright n=\dots=a_{k}\upharpoonright n$, we have that
$[M_{i}]_{i<\omega}\models R(a_{1},\dots,a_{k})\iff M_{n}\models R(a_{1}(n),\dots,R(a_{k}(n)).$
While most of the results in the following can be stated more generally (see [21, Definition 4.5]), for our purposes, we restrict our attention to products as in Definition 7.9.
Throughout this section, we fix a family $(M_{i})_{i\in\omega}$ of Fraïssé structures, where each $M_{i}$ is transitive and has domain $\omega$. (This is called $\omega$-pure in [21, Definition 5.2].)
We fix a family $\{s_{n}\mid n<\omega\}$ of binary relation symbols, and let $M_{i}^{s}$ be the expansion of $M_{i}$ by $\{s_{n}\mid n<\omega\}$ such that $\left(s_{n}\right)^{M_{i}^{s}}=\left\{\begin{array}[]{ll}\left(M_{i}\right)^{2}&\mbox{if }i>n\\ \varnothing&\mbox{if }i\leq n.\end{array}\right.$.
Remark 7.10 ([21, Remark 5.4]).
With the expansion above, $[M_{i}^{s}]_{i<\omega}$ has refining equivalence relations $s_{0}\supseteq s_{1}\supseteq s_{2}\supseteq\cdots$. We also have that $[M_{i}^{s}]_{i<\omega}$ is the expansion of $[M_{i}]_{i<\omega}$ by $\{s_{n}|n<\omega\}$ such that $s_{n}(x,y)\Leftrightarrow x(i)=y(i)$ for all $i\leq n$.
The following lemma is a special case of [21, Lemma 4.15], and its proof is straightforward.
@@
For $j\geq 1$, let $N_{j}=M_{j}$. Let $N^{s}=[N_{j}^{s}]_{j\geq 1}$ (here the enumeration of the $s_{n}$ begins at $n=1$). Then we have $M_{0}[N^{s}]^{s_{0}}\cong[M_{i}^{s}]_{i<\omega}$ via the map $(a,b)\mapsto a\smallfrown~{}b$ (the concatenation of $a$ and $b$).
@@
A substructure $D\subseteq[M^{s}_{i}]_{i<\omega}$ is dense if for all $t\in\omega^{<\omega}$, there is some $d\in D$ such that $d\supseteq t$.
Fact 7.13 ([21, Theorem 5.14]).
@@
    Up to isomorphism, there is a unique countable dense substructure $D\subseteq[M_{i}^{s}]_{i<\omega}$.
@@
    Such a $D$ is transitive and ultrahomogeneous.
Fact 7.14 ([21, Theorem 6.1]).
If $D\subseteq[M_{i}^{s}]_{i<\omega}$ is a countable dense substructure, then $D$ is indivisible.
@@
Let $D\subseteq[M_{i}^{s}]_{i<\omega}$ be a countable dense substructure. Then the following are equivalent:
@@
Let $D\subseteq[M_{i}^{s}]_{i<\omega}$ and $D^{\prime}\subseteq N^{s}$ be countable dense substructures. Then $D\cong M_{0}[D^{\prime}]^{s_{0}}$.
@@
As $M_{0}[D^{\prime}]^{s_{0}}$ is countable and dense in $M_{0}[N^{s}]^{s_{0}}$, by Lemma 7.11 and Fact 7.13 we have $D\cong M_{0}[D^{\prime}]^{s_{0}}$. ∎
@@
Let $D\subseteq[M_{i}^{s}]_{i<\omega}$ be a countable dense substructure. Suppose that $M_{0}$ has FPTn. Then $D$ has FPTn.
@@
Let $D^{\prime}\subseteq N^{s}$ be countable and dense. By Lemma 7.16, $D\cong M_{0}[D^{\prime}]^{s_{0}}$. By Proposition 7.8, as $M_{0}$ has FPTn, we have that $M_{0}[D^{\prime}]^{s_{0}}$ has FPTn. ∎
@@
By Proposition 7.17, $D$ has FPT if and only if $M_{0}$ has FPT. By Fact 7.13, $D$ is ultrahomogeneous. Recall that $M_{0}$ was also assumed to be ultrahomogeneous and countable. Therefore, by Theorem 5.1, we have that $D$ has the externally definable Ramsey property if and only if $D$ has FPT, and $M_{0}$ has the externally definable Ramsey property if and only if $M_{0}$ has FPT. ∎
@@
If $M_{i}$ is an infinite set in the language of pure equality, for all $i\in\omega$, then a countable dense substructure $D\subseteq[M_{i}^{s}]_{i<\omega}$ is the structure with infinitely many refining equivalence relations (unsaturated version) defined in Section 6.8.
@@
The partial type $\{x\neq y\}\,\cup\,\{s_{i}(x,y)\mid i\in\omega\}$ is finitely satisfiable but not satisfiable in $D$. ∎
@@
Let $\mathcal{L}$ be a countable relational language, and let $T$ be a universal $\mathcal{L}$-theory closed under lexicographic products, i.e. if $M,N\models T$ then $M[N]\models T$. Let $T^{s}$ be the $\mathcal{L}\cup\{s_{i}\mid i\in\omega\}$-theory defined as $T$ together with the $\{s_{i}\mid i\in\omega\}$-theory of infinitely many refining equivalence relations with infinite classes.
@@
Take $A_{0},A_{1}$ to be two distinct countable transitive ultrahomogeneous models of $T$, such that $A_{0}$ has the externally definable RP. For each $a\in 2^{\omega\setminus\{0\}}$, let $M(a)_{0}=A_{0}$ and $M(a)_{j}=A_{a(j)}$ for $j\geq 1$. Let $M(a)$ be a countable dense substructure of $[M(a)_{i}^{s}]_{i<\omega}$. By Fact 7.13, $M(a)$ is ultrahomogeneous. By Theorem 7.15, $M(a)$ has the externally definable Ramsey property. By Proposition 7.19, $M(a)$ is not $\aleph_{0}$-saturated.
It remains to show that $M(a)\not\cong M(b)$ for distinct $a,b\in 2^{\omega\setminus\{0\}}$. Assume for a contradiction that $f:M(a)\to M(b)$ is an isomorphism. Let $m<\omega$ be least such that $a(m+1)\neq b(m+1)$. By density of $M(a)$, for each $i<\omega$ there is $\sigma_{i}\in M(a)$ with initial segment $(a(0),\cdots,a(m),i)$. Notice that the $\sigma_{i}$ are pairwise $s_{m}$-equivalent, but not $s_{m+1}$-equivalent, and, in addition, they $s_{m+1}$-cover their equivalence class. As $f$ is an $\mathcal{L}\cup\{s_{i}\mid i<\omega\}$-isomorphism, the same properties are also satisfied by the $f(\sigma_{i})$ in $M(b)$. Finally, notice that, by the definition of the infinite lexicographic product, the substructure induced on $\{\sigma_{i}\}_{i\in\omega}$ in $M(a)$ is isomorphic to $M(a)_{m+1}$ and the substructure induced on $\{f(\sigma_{i})\}_{i\in\omega}$ in $M(b)$ is isomorphic to $M(b)_{m+1}$. But $M(b)_{m+1}\cong A_{b(m+1)}\not\cong A_{a(m+1)}\cong M(a)_{m+1}$, contradiction. ∎
@@
  • the theory of $n$-hypergraphs, for $n\geq 2$;
  • the theory of $K_{m}$-free $n$-hypergraphs for $m\geq n\geq 3$;
  • the theory of partially ordered sets.
@@ 8. Further questions
Does there exist an ultrahomogeneous structure $M$ with uncountable age such that $M$ has the externally definable Ramsey property but does not have FPT?
@@
In this paper, for most examples given of Fraïssé structures $M$ with the externally definable Ramsey property, we show this property by checking the $\bar{a}$-RP for all colourings, for $|\bar{a}|<t$ (where $t$ is the maximum arity of $\mathcal{L}$ and we have QE). Can we find interesting examples of Fraïssé structures $M$ where $M$ has the externally definable Ramsey property, but does not have the $\bar{a}$-RP for $|\bar{a}|<t$? (See Section 6.2.)
@@
We would also like to know the following regarding the fixed points on type spaces property (FPT):
Question 8.11.
Do there exist Fraïssé structures $M,M^{\prime}$ (possibly in different languages), with $\operatorname{Aut}(M),\operatorname{Aut}(M^{\prime})$ isomorphic as topological groups, such that $M$ has FPT and $M^{\prime}$ does not?